For a particular double-slit experiment, the slit separation is 4.50 μm, the screen is located 4.00 m from the slits, and the wavelength of the light illuminating the slits is 580 nm (yellow). You have a detector located on the screen at a distance of 20.5 cm from the center of the interference pattern. (a) Where in the pattern is this point located – at a minimum, maximum or in-between? Give the m values associated with this position (e.g. at the m = 3 maximum, between the 3rd minimum and m = 4 maximum, etc.). (b) What is the ratio of the intensity at this point to that at the center of the interference pattern?
Answer #1

d=4.50um Ys nodanx sporiox 4 - 430x106 m = nx o. 5155 - ni 51155 am I n=0,1,2.. I moro Inn = (MAY) & D = (n/a) 51.55 - n5o,